OpenStudy (anonymous):

given the following revenue and cost functions, find the x-value that makes profit a maximum: R(x)=57x-2x^2 C(x)=21x+106 (remember that profit equals revenue minus cost)

OpenStudy (anonymous):

First step: Subtract the cost function from the revenue function to give a profit function. Then find the maximum of the profit function.

OpenStudy (anonymous):

would it look like this: (f(x)=57x-2x^2)-(f(x)=21x+106)

OpenStudy (anonymous):

\(R(x)=57x-2x^2\) \(C(x)=21x+106\) \(P(x)=R(x)-C(x)\) \(P(x)=57x-2x^2-(21x + 106)\) \(P(x)=-2x^2+36x-106\)

OpenStudy (anonymous):

k, give me a minute to digest...

OpenStudy (anonymous):

so then P(9)=54?

OpenStudy (anonymous):

P(9)=56, and 9 is the correct answer, yes.

OpenStudy (anonymous):

thank you.